# Chirped amplitude mode in photo-excited superconductors

### v1

@@ Appendix A Implementation details
$\bar{t}_{0}$ $=\begin{pmatrix}e^{i\phi(t)}&&0\\ 0&&e^{-i\phi(t)}\end{pmatrix},$ (13)
@@
due to the Peierls substitution.
@@
We use the fifth-order backward differentiation formula multistep method to discretize the KBE. Gregory quadrature is used to discretize memory integrals to fifth-order accuracy. The discrete Lehmann representation [72], implemented in libdlr [73], is used to discretize the imaginary time variables appearing in the vertical leg of the Keldysh contour, as in Ref. [74]. We have confirmed that the compression of the Green’s function using the HODLR scheme preserves conservation laws, namely the density and energy, to within the SVD truncation tolerance used in the scheme, which we set to $10^{-8}$.
@@
Appendix B Numerical ranks in HODLR scheme
Figure 5: (a) Rank of the truncated SVD of the largest block in a HODLR decompostion of Green’s function at timestep $T$. Results for several different pump strengths are presented. $\beta=6$ is a result in the equilibrium disordered state. (b) Total memory required to store Green’s function and self-energy for the presented pump strengths.

### v2

@@ Appendix A Implementation details
$\bar{t}_{0}$ $=\begin{pmatrix}e^{iA(t)}&&0\\ 0&&e^{-iA(t)}\end{pmatrix},$ (9)
@@
due to the Peierls substitution. We set the charge and the lattice distance to unity, $q=a=1$.
@@
We use the fifth-order backward differentiation formula multistep method to discretize the KBE. Gregory quadrature is used to discretize memory integrals to fifth-order accuracy. The discrete Lehmann representation [83], implemented in libdlr [84], is used to discretize the imaginary time variables appearing in the vertical leg of the Keldysh contour, as in Ref. [85]. We have confirmed that the compression of the Green’s function using the HODLR scheme preserves conservation laws, namely the density and energy, to within the Singular Value Decomposition truncation tolerance used in the scheme, which we set to $10^{-8}$.
@@
Appendix B Hartree-Fock
Solving the KBE allows for inclusion of dynamical fluctuations absent in a Hartree-Fock treatment of superconductors. We show that including these correlations leads to qualitatively different dynamics, demonstrating the necessity of calculations beyond the mean field level. In the main text, we observed chirping of the amplitude mode after the photo-excitation and we contrast these dynamics with the Hartree-Fock approximation, see Fig. 5. Already in equilibrium, the critical inverse temperature for the transition is $\beta_{c}=2$ which is much higher than in the 2nd Born approximation.
Figure 5: Time evolution of the order parameter $\phi$ within the Hartree-Fock approximation for increasing strength of the external pulse $E_{0}$.
We can use this order parameter data to confirm one of the main predictions of a Hartree-Fock treatment, which is a power-law decay of the Higgs mode [25]. The amplitude of the Higgs oscillation in the Hartree-Fock and second Born approximation is presented in Fig. 6. In the Hartree-Fock case, the expected power-law decay of $\phi\propto t^{-1/2}$ is shown to hold true over two orders of magnitude. However, when fluctuations are introduced through the 2nd Born self-energy, the Higgs mode instead decays exponentially [26], with the lifetime strongly depending on the excitation strength.
Figure 6: (a) Decay of the oscillating amplitude of the order parameter, $A$, for different excitation strengths $E_{0}$ within the 2nd Born  (a) and the Hartree-Fock approximation (c). The amplitude of the Hartree-Fock results is compared with the power-law decay $t^{-1/2}$. (b) The initial amplitude, $A_{0}$, and inverse lifetime, $1/T_{L}$, of the Higgs mode with increasing pump intensity for the 2B results.
Fig. 5 shows that when the pump field is turned off, the order parameter oscillates around a fixed value, in stark contrast to the continual decay present in the second order theory. This suggests an immediate alteration of the potential energy surface within the system, suggesting an absence of chirping. This absence of chirping is confirmed in Fig. 7, where the time dependence of the Higgs oscillation frequency is presented. The frequency is constant for all pump strengths, with only noise less than 1% of the original frequency.
Figure 7: Windowed Fourier transform $F[\phi](t_{w})$ of the order parameter centered around $t_{w}$ with background subtracted from the Hartree-Fock dynamics. Results for pump amplitudes (a) $E_{0}=0.6$, (b) $E_{0}=0.5$, (c) $E_{0}=0.4$, (d) $E_{0}=0.3$. The red line tracks the maximum of the spectrum, and varies by at most one unit of the frequency resolution.
Appendix C Numerical ranks in HODLR scheme
Figure 8: (a) Rank of the truncated SVD of the largest block in a HODLR decompostion of Green’s function at timestep $T$. Results for several different pump strengths are presented. (b) Compression factor for Green’s function and self-energy for the presented pump strengths.